OpenStudy (anonymous):

Look and see if I am right... 1. Suppose the function f of x equals 4 divided by x a) Use the definition of the derivative to show that f ' (-2) = -1. b) Write an equation for the line tangent to the graph of f at x = -2. f(x) = 4/x = 4*x^(-1) a. f ' (x) = -4*x^(-2) f'(x) = -4*(-2)^(-2)=-1

OpenStudy (anonymous):

b) At x = -2, f ' (-2) = slope : y - y_1 = slope * (x - x_1) where [x_1, y_1] = [-2, f(-2)]

OpenStudy (kc_kennylau):

Therefore you plug in the values: \[y-f(-2)=-1\times(x-(-2))\]

OpenStudy (anonymous):

For which part?

OpenStudy (kc_kennylau):

for part b

OpenStudy (anonymous):

okay did part a look alright?

OpenStudy (kc_kennylau):

yep :D

OpenStudy (anonymous):

okay so what do I plug into the second one?

OpenStudy (kc_kennylau):

what do you mean the second one

OpenStudy (anonymous):

part b

OpenStudy (kc_kennylau):

you plug in (x_1,y_1)=(-2,f(-2)) to y-y_1=slope(x-x_1)

OpenStudy (kc_kennylau):

That gives you: \[y-f(-2)=-1\times(x-(-2))\]

OpenStudy (anonymous):

and that is all I have to do?

OpenStudy (kc_kennylau):

ofc :)

OpenStudy (kc_kennylau):

well you need to simplify it

OpenStudy (anonymous):

-x+2?

OpenStudy (kc_kennylau):

y-f(-2)=-(x+2) y-4/(-2)=-x-2 y+2=-x-2 x+y+4=0

OpenStudy (anonymous):

sorry i meant -x-2

OpenStudy (anonymous):

okay awesome. Thank you!